For the reaction H2(g) + F2(g) ↔ 2 HF(g), the value of Kc = 195 at some temperature. The initial conditions are 0.50HF and none of the LHS materials. Decide in which direction the reaction will spontaneously change. Then use the ICE scheme, set up properly, to solve for “x.” Once x is determined, fill in the bottom line to find all three equilibrium concentrations.


1
Expert's answer
2020-07-14T15:16:20-0400

 H2(g) + F2(g) ↔ 2 HF(g), the value of Kc = 195 at some temperature


So 2 HF ↔ H2 +F2 , the value of Kc, = (1/195 )at some temperature .

initial 0.50M 0 0

equillibrium 0.5(1- x) 0.5x 0.5x


 Kc, = 0.5x * 0.5x / 0.5 ( 1-x) = 1 / 195 .


1- x = 1 ( approx ) .


( 0.5x )2 = ( 0.5 ) * ( 1/295 ) = 0.001694


0.5x = [H2] = [F2] = 0.041158 .



FORMATION OF HF FROM H2 AND F2 ARE MORE favorable THEN dissociation OF HF TO H2 AND F2 .